OpenStudy (anonymous):

PLEASE HELP! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

y = -6x - 15 -3 = -6(-2) - 15 -3 = 12 - 15 -3 = -3 So the first one.